a bicycle tire has a diameter of 2 feet. How many revolutions will it make in going a distance of 50 feet?
Answer by Alan3354(69443)

a bicycle tire has a diameter of 2 feet. How many revolutions will it make in going a distance of 50 feet?
-----------
It moves 2pi feet per rev
# of revs = 50/2pi = 25/pi